# Tracing Galaxy Evolution in the Nearby Universe: The Role of Dark Matter

### v1

@@ Appendix B Stellar mass vs redshift
Figure B1: Behavior of stellar mass $({\bf M_{EN}})$ as function of redshift for quasi-constant dynamical mass. Each colour and symbol represents quasi-constant dynamical mass. The lower-left part of the graph (black dots) corresponds to $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ $\sim$ 10.30 while the upper-right part of the graph (blue triangles) corresponds to $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ $\sim$ 11.80. The difference in $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ between consecutive symbols is approximately 0.5. The mean uncertainty of the $\log({\bf M_{EN}/{\bf M_{Solar}}})$ is approximately 0.055 dex.
@@
Figure B2: Behavior of stellar mass $({\bf M_{WD}})$ as function of redshift for quasi-constant dynamical mass. Each colour and symbol represents quasi-constant dynamical mass. The lower-left part of the graph (black dots) corresponds to $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ $\sim$ 10.30 while the upper-right part of the graph (blue triangles) corresponds to $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ $\sim$ 11.80. The difference in $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ between consecutive symbols is approximately 0.5. The mean error of the $\log({\bf M_{WD}/{\bf M_{Solar}}})$ is approximately 0.092 dex.
@@
Figure B3: Behavior of stellar mass $({\bf M_{WN}})$ as function of redshift for quasi-constant dynamical mass. Each colour and symbol represents quasi-constant dynamical mass. The lower-left part of the graph (black dots) corresponds to $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ $\sim$ 10.30 while the upper-right part of the graph (blue triangles) corresponds to $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ $\sim$ 11.80. The difference in $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ between consecutive symbols is approximately 0.5. The mean uncertainty of the $\log({\bf M_{WN}/{\bf M_{Solar}}})$ is approximately 0.070 dex.
@@
Figure B4: Behavior of stellar mass $({\bf M_{PP}})$ as function of redshift for quasi-constant dynamical mass. Each colour and symbol represents quasi-constant dynamical mass. The lower-left part of the graph (black dots) corresponds to $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ $\sim$ 10.30 while the upper-right part of the graph (blue triangles) corresponds to $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ $\sim$ 11.80. The difference in $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ between consecutive symbols is approximately 0.5. The mean error of the $\log({\bf M_{PP}/{\bf M_{Solar}}})$ is approximately 0.089 dex.
@@
Figure B5: Behavior of stellar mass $({\bf M_{P03}})$ as function of redshift for quasi-constant dynamical mass. Each colour and symbol represents quasi-constant dynamical mass. The lower-left part of the graph (black dots) corresponds to $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ $\sim$ 10.30 while the upper-right part of the graph (blue triangles) corresponds to $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ $\sim$ 11.80. The difference in $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ between consecutive symbols is approximately 0.5. The mean uncertainty of the $\log({\bf M_{P03}/{\bf M_{Solar}}})$ is approximately 0.214 dex.
@@
Figure B6: Behavior of stellar mass $({\bf M_{P11}})$ as function of redshift for quasi-constant dynamical mass. Each colour and symbol represents quasi-constant dynamical mass. The lower-left part of the graph (black dots) corresponds to $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ $\sim$ 10.30 while the upper-right part of the graph (blue triangles) corresponds to $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ $\sim$ 11.80. The difference in $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ between consecutive symbols is approximately 0.5. The mean error of the $\log({\bf M_{P11}/{\bf M_{Solar}}})$ is approximately 0.174 dex.
@@
Figure B7: Behavior of stellar mass $({\bf M_{SF}})$ as function of redshift for quasi-constant dynamical mass. Each colour and symbol represents quasi-constant dynamical mass. The lower-left part of the graph (black dots) corresponds to $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ $\sim$ 10.30 while the upper-right part of the graph (blue triangles) corresponds to $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ $\sim$ 11.80. The difference in $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ between consecutive symbols is approximately 0.5. The mean uncertainty of the $\log({\bf M_{SF}/{\bf M_{Solar}}})$ is approximately 0.128 dex.
Appendix C Dynamical mass - stellar mass vs dynamical mass
@@
Figure C1: Difference between dynamical and stellar mass $({\bf M_{EN}})$ as function of dynamical mass for the LTGs samples. Each colour and symbol represents a quasi-constant redshift. The upper-left part of the graph (black dots) corresponds to $z\sim 0.025$ while the lower-right part of the graph (dark green triangles) corresponds to $z\sim 0.165$. The difference in redshift between consecutive symbols is approximately 0.01. The mean uncertainty of the difference between $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ and $\log({\bf M_{EN}/{\bf M_{Solar}}})$ is approximately 0.279 dex. The left column of the mosaic contains the full range of redshift, while the right column only contains three specific redshifts: low (black dots $z\sim 0.025$), intermediate (purple squares $z\sim 0.095$), and high (dark green triangles $z\sim 0.165$), the latter with the aim of appreciating more clearly the differences in dynamical and stellar masses due to redshift and also to display the uncertainties for each point.
@@
Figure C2: Difference between dynamical and stellar mass $({\bf M_{WD}})$ as function of dynamical mass for the LTGs samples. Each colour and symbol represents a quasi-constant redshift. The upper-left part of the graph (black dots) corresponds to $z\sim 0.025$ while the lower-right part of the graph (dark green triangles) corresponds to $z\sim 0.165$. The difference in redshift between consecutive symbols is approximately 0.01. The mean uncertainty of the difference between $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ and $\log({\bf M_{WD}/{\bf M_{Solar}}})$ is approximately 0.285 dex. The left column of the mosaic contains the full range of redshift, while the right column only contains three specific redshifts: low (black dots $z\sim 0.025$), intermediate (purple squares $z\sim 0.095$), and high (dark green triangles $z\sim 0.165$), the latter with the aim of appreciating more clearly the differences in dynamical and stellar masses due to redshift and also to display the uncertainties for each point.
@@
Figure C3: Difference between dynamical and stellar mass $({\bf M_{WN}})$ as function of dynamical mass for the LTGs samples. Each colour and symbol represents a quasi-constant redshift. The upper-left part of the graph (black dots) corresponds to $z\sim 0.025$ while the lower-right part of the graph (dark green triangles) corresponds to $z\sim 0.165$. The difference in redshift between consecutive symbols is approximately 0.01. The mean uncertainty of the difference between $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ and $\log({\bf M_{WN}/{\bf M_{Solar}}})$ is approximately 0.281 dex. The left column of the mosaic contains the full range of redshift, while the right column only contains three specific redshifts: low (black dots $z\sim 0.025$), intermediate (purple squares $z\sim 0.095$), and high (dark green triangles $z\sim 0.165$), the latter with the aim of appreciating more clearly the differences in dynamical and stellar masses due to redshift and also to display the uncertainties for each point.
@@
Figure C4: Difference between dynamical and stellar mass $({\bf M_{PP}})$as function of dynamical mass for the LTGs samples. Each colour and symbol represents a quasi-constant redshift. The upper-left part of the graph (black dots) corresponds to $z\sim 0.025$ while the lower-right part of the graph (dark green triangles) corresponds to $z\sim 0.165$. The difference in redshift between consecutive symbols is approximately 0.01. The mean uncertainty of the difference between $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ and $\log({\bf M_{PP}/{\bf M_{Solar}}})$ is approximately 0.284 dex. The left column of the mosaic contains the full range of redshift, while the right column only contains three specific redshifts: low (black dots $z\sim 0.025$), intermediate (purple squares $z\sim 0.095$), and high (dark green triangles $z\sim 0.165$), the latter with the aim of appreciating more clearly the differences in dynamical and stellar masses due to redshift and also to display the uncertainties for each point.
@@
Figure C5: Difference between dynamical and stellar mass $({\bf M_{P03}})$as function of dynamical mass for the LTGs samples. Each colour and symbol represents a quasi-constant redshift. The upper-left part of the graph (black dots) corresponds to $z\sim 0.025$ while the lower-right part of the graph (dark green triangles) corresponds to $z\sim 0.165$. The difference in redshift between consecutive symbols is approximately 0.01. The mean uncertainty of the difference between $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ and $\log({\bf M_{P03}/{\bf M_{Solar}}})$ is approximately 0.331 dex. The left column of the mosaic contains the full range of redshift, while the right column only contains three specific redshifts: low (black dots $z\sim 0.025$), intermediate (purple squares $z\sim 0.095$), and high (dark green triangles $z\sim 0.165$), the latter with the aim of appreciating more clearly the differences in dynamical and stellar masses due to redshift and also to display the uncertainties for each point.
@@
Figure C6: Difference between dynamical and stellar mass $({\bf M_{P11}})$as function of dynamical mass for the LTGs samples. Each colour and symbol represents a quasi-constant redshift. The upper-left part of the graph (black dots) corresponds to $z\sim 0.025$ while the lower-right part of the graph (green triangles) corresponds to $z\sim 0.165$. The difference in redshift between consecutive symbols is approximately 0.01. The mean uncertainty of the difference between $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ and $\log({\bf M_{P11}/{\bf M_{Solar}}})$ is approximately 0.312 dex. The left column of the mosaic contains the full range of redshift, while the right column only contains three specific redshifts: low (black dots $z\sim 0.025$), intermediate (purple squares $z\sim 0.095$), and high (dark green triangles $z\sim 0.165$), the latter with the aim of appreciating more clearly the differences in dynamical and stellar masses due to redshift and also to display the uncertainties for each point.
@@
Figure C7: Difference between dynamical and stellar mass $({\bf M_{SF}})$as function of dynamical mass for the LTGs samples. Each colour and symbol represents a quasi-constant redshift. The upper-left part of the graph (black dots) corresponds to $z\sim 0.025$ while the lower-right part of the graph (green triangles) corresponds to $z\sim 0.165$. The difference in redshift between consecutive symbols is approximately 0.01. The mean uncertainty of the difference between $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ and $\log({\bf M_{SF}/{\bf M_{Solar}}})$ is approximately 0.294 dex. The left column of the mosaic contains the full range of redshift, while the right column only contains three specific redshifts: low (black dots $z\sim 0.025$), intermediate (purple squares $z\sim 0.095$), and high (dark green triangles $z\sim 0.165$), the latter with the aim of appreciating more clearly the differences in dynamical and stellar masses due to redshift and also to display the uncertainties for each point.

### v2

@@ Appendix B Stellar mass versus redshift
Figure B1: Behavior of stellar mass $({\bf M_{EN}})$ as function of the redshift for quasi-constant dynamical mass. Each color and symbol represents quasi-constant dynamical mass. The lower-left part of the graph (black dots) corresponds to $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ $\sim$ 10.30 while the upper-right part of the graph (blue triangles) corresponds to $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ $\sim$ 11.80. The difference in $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ between consecutive symbols is approximately 0.5. The mean uncertainty of the $\log({\bf M_{EN}/{\bf M_{Solar}}})$ is approximately 0.055 dex.
@@
Figure B2: Behavior of stellar mass $({\bf M_{WD}})$ as function of the redshift for quasi-constant dynamical mass. Each color and symbol represents quasi-constant dynamical mass. The lower-left part of the graph (black dots) corresponds to $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ $\sim$ 10.30 while the upper-right part of the graph (blue triangles) corresponds to $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ $\sim$ 11.80. The difference in $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ between consecutive symbols is approximately 0.5. The mean error of the $\log({\bf M_{WD}/{\bf M_{Solar}}})$ is approximately 0.092 dex.
@@
Figure B3: Behavior of stellar mass $({\bf M_{WN}})$ as function of the redshift for quasi-constant dynamical mass. Each color and symbol represents quasi-constant dynamical mass. The lower-left part of the graph (black dots) corresponds to $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ $\sim$ 10.30 while the upper-right part of the graph (blue triangles) corresponds to $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ $\sim$ 11.80. The difference in $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ between consecutive symbols is approximately 0.5. The mean uncertainty of the $\log({\bf M_{WN}/{\bf M_{Solar}}})$ is approximately 0.070 dex.
@@
Figure B4: Behavior of stellar mass $({\bf M_{PP}})$ as function of the redshift for quasi-constant dynamical mass. Each color and symbol represents quasi-constant dynamical mass. The lower-left part of the graph (black dots) corresponds to $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ $\sim$ 10.30 while the upper-right part of the graph (blue triangles) corresponds to $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ $\sim$ 11.80. The difference in $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ between consecutive symbols is approximately 0.5. The mean error of the $\log({\bf M_{PP}/{\bf M_{Solar}}})$ is approximately 0.089 dex.
@@
Figure B5: Behavior of stellar mass $({\bf M_{P03}})$ as function of the redshift for quasi-constant dynamical mass. Each color and symbol represents quasi-constant dynamical mass. The lower-left part of the graph (black dots) corresponds to $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ $\sim$ 10.30 while the upper-right part of the graph (blue triangles) corresponds to $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ $\sim$ 11.80. The difference in $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ between consecutive symbols is approximately 0.5. The mean uncertainty of the $\log({\bf M_{P03}/{\bf M_{Solar}}})$ is approximately 0.214 dex.
@@
Figure B6: Behavior of stellar mass $({\bf M_{P11}})$ as function of the redshift for quasi-constant dynamical mass. Each color and symbol represents quasi-constant dynamical mass. The lower-left part of the graph (black dots) corresponds to $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ $\sim$ 10.30 while the upper-right part of the graph (blue triangles) corresponds to $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ $\sim$ 11.80. The difference in $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ between consecutive symbols is approximately 0.5. The mean error of the $\log({\bf M_{P11}/{\bf M_{Solar}}})$ is approximately 0.174 dex.
@@
Figure B7: Behavior of stellar mass $({\bf M_{SF}})$ as function of the redshift for quasi-constant dynamical mass. Each color and symbol represents quasi-constant dynamical mass. The lower-left part of the graph (black dots) corresponds to $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ $\sim$ 10.30 while the upper-right part of the graph (blue triangles) corresponds to $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ $\sim$ 11.80. The difference in $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ between consecutive symbols is approximately 0.5. The mean uncertainty of the $\log({\bf M_{SF}/{\bf M_{Solar}}})$ is approximately 0.128 dex.
Appendix C Dynamical mass - stellar mass versus dynamical mass
This appendix contains the complete set of mosaics for the logarithmic mass difference between dynamical and stellar mass ($\Delta\log\mathbf{M}$) as a function of the dynamical mass. These figures demonstrate the consistency of the observed evolutionary trends across different mass estimation models (see Figures C1C7).
@@ Appendix C Dynamical mass - stellar mass versus dynamical mass
Figure C1: Difference between dynamical and stellar mass $({\bf M_{EN}})$ as function of the dynamical mass for the LTGs samples. Each color and symbol represents a quasi-constant redshift. The upper-left part of the graph (black dots) corresponds to $z\sim 0.025$ while the lower-right part of the graph (dark green triangles) corresponds to $z\sim 0.165$. The difference in redshift between consecutive symbols is approximately 0.01. The mean uncertainty of the difference between $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ and $\log({\bf M_{EN}/{\bf M_{Solar}}})$ is approximately 0.279 dex. The left column of the mosaic contains the full range of redshift, while the right column only contains three specific redshifts: low (black dots, $z\sim 0.025$), intermediate (purple squares, $z\sim 0.095$), and high (dark green triangles, $z\sim 0.165$), the latter with the aim of appreciating more clearly the differences in dynamical and stellar masses due to redshift and also to display the uncertainties for each point.
@@
Figure C2: Difference between dynamical and stellar mass $({\bf M_{WD}})$ as function of dynamical mass for the LTGs samples. Each color and symbol represents a quasi-constant redshift. The upper-left part of the graph (black dots) corresponds to $z\sim 0.025$ while the lower-right part of the graph (dark green triangles) corresponds to $z\sim 0.165$. The difference in redshift between consecutive symbols is approximately 0.01. The mean uncertainty of the difference between $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ and $\log({\bf M_{WD}/{\bf M_{Solar}}})$ is approximately 0.285 dex. The left column of the mosaic contains the full range of redshift, while the right column only contains three specific redshifts: low (black dots $z\sim 0.025$), intermediate (purple squares $z\sim 0.095$), and high (dark green triangles $z\sim 0.165$), the latter with the aim of appreciating more clearly the differences in dynamical and stellar masses due to redshift and also to display the uncertainties for each point.
@@
Figure C3: Difference between dynamical and stellar mass $({\bf M_{WN}})$ as function of dynamical mass for the LTGs samples. Each color and symbol represents a quasi-constant redshift. The upper-left part of the graph (black dots) corresponds to $z\sim 0.025$ while the lower-right part of the graph (dark green triangles) corresponds to $z\sim 0.165$. The difference in redshift between consecutive symbols is approximately 0.01. The mean uncertainty of the difference between $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ and $\log({\bf M_{WN}/{\bf M_{Solar}}})$ is approximately 0.281 dex. The left column of the mosaic contains the full range of redshift, while the right column only contains three specific redshifts: low (black dots $z\sim 0.025$), intermediate (purple squares $z\sim 0.095$), and high (dark green triangles $z\sim 0.165$), the latter with the aim of appreciating more clearly the differences in dynamical and stellar masses due to redshift and also to display the uncertainties for each point.
@@
Figure C4: Difference between dynamical and stellar mass $({\bf M_{PP}})$as function of dynamical mass for the LTGs samples. Each color and symbol represents a quasi-constant redshift. The upper-left part of the graph (black dots) corresponds to $z\sim 0.025$ while the lower-right part of the graph (dark green triangles) corresponds to $z\sim 0.165$. The difference in redshift between consecutive symbols is approximately 0.01. The mean uncertainty of the difference between $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ and $\log({\bf M_{PP}/{\bf M_{Solar}}})$ is approximately 0.284 dex. The left column of the mosaic contains the full range of redshift, while the right column only contains three specific redshifts: low (black dots $z\sim 0.025$), intermediate (purple squares $z\sim 0.095$), and high (dark green triangles $z\sim 0.165$), the latter with the aim of appreciating more clearly the differences in dynamical and stellar masses due to redshift and also to display the uncertainties for each point.
@@
Figure C5: Difference between dynamical and stellar mass $({\bf M_{P03}})$as function of dynamical mass for the LTGs samples. Each color and symbol represents a quasi-constant redshift. The upper-left part of the graph (black dots) corresponds to $z\sim 0.025$ while the lower-right part of the graph (dark green triangles) corresponds to $z\sim 0.165$. The difference in redshift between consecutive symbols is approximately 0.01. The mean uncertainty of the difference between $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ and $\log({\bf M_{P03}/{\bf M_{Solar}}})$ is approximately 0.331 dex. The left column of the mosaic contains the full range of redshift, while the right column only contains three specific redshifts: low (black dots $z\sim 0.025$), intermediate (purple squares $z\sim 0.095$), and high (dark green triangles $z\sim 0.165$), the latter with the aim of appreciating more clearly the differences in dynamical and stellar masses due to redshift and also to display the uncertainties for each point.
@@
Figure C6: Difference between dynamical and stellar mass $({\bf M_{P11}})$as function of dynamical mass for the LTGs samples. Each color and symbol represents a quasi-constant redshift. The upper-left part of the graph (black dots) corresponds to $z\sim 0.025$ while the lower-right part of the graph (green triangles) corresponds to $z\sim 0.165$. The difference in redshift between consecutive symbols is approximately 0.01. The mean uncertainty of the difference between $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ and $\log({\bf M_{P11}/{\bf M_{Solar}}})$ is approximately 0.312 dex. The left column of the mosaic contains the full range of redshift, while the right column only contains three specific redshifts: low (black dots $z\sim 0.025$), intermediate (purple squares $z\sim 0.095$), and high (dark green triangles $z\sim 0.165$), the latter with the aim of appreciating more clearly the differences in dynamical and stellar masses due to redshift and also to display the uncertainties for each point.
@@
Figure C7: Difference between dynamical and stellar mass $({\bf M_{SF}})$as function of dynamical mass for the LTGs samples. Each color and symbol represents a quasi-constant redshift. The upper-left part of the graph (black dots) corresponds to $z\sim 0.025$ while the lower-right part of the graph (green triangles) corresponds to $z\sim 0.165$. The difference in redshift between consecutive symbols is approximately 0.01. The mean uncertainty of the difference between $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ and $\log({\bf M_{SF}/{\bf M_{Solar}}})$ is approximately 0.294 dex. The left column of the mosaic contains the full range of redshift, while the right column only contains three specific redshifts: low (black dots $z\sim 0.025$), intermediate (purple squares $z\sim 0.095$), and high (dark green triangles $z\sim 0.165$), the latter with the aim of appreciating more clearly the differences in dynamical and stellar masses due to redshift and also to display the uncertainties for each point.